If the area is 144sqcm & Perimeter is 48
Answer by JBarnum(2146)

This is not a question but just at glancing you should make sure u memorize your squares you know like 5*5=25 7*7=49 9*9=81 11*11=121 12*12=144 also know that a rectangle has 4 sides and since this could be a square you should know your multiples of basic numbers like 2 4 6 8 or 5 10 15 20 or 12 24 36 48. 48/4=12 144/12=12